OpenStudy (anonymous):

A fund drive announces that it has rasied $8,032, which is 16% of its goal. Q: How much more must be raised?

OpenStudy (anonymous):

8032/x=16/100 16x=803200 x=$50200

OpenStudy (anonymous):

8032 = .16x Using inverse operations, divide each side by .16. This will make x = 50200 total. What you really want to find out is the 84%, so you will take 50200 and subtract 8032 to $42,168, which is the answer to your question.